OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2.

OpenStudy (trojanpoem):

-3 - 6 / (10-x) = -3/2 -9/10-x = -3/2 -9 * 2 = -3 (10-x) -18 = -30 + 3x 3x = 12 , x= 4